Please answer the "Required 1" showing the apportionment of cost, cost of each machine, depreciation or amortization expense, journal entry, and journal entries for each machine: Compute the cost of each machine by making a supportable allocation of the total cost to the three machines (round your calculations to two decimal places). Explain the rationale for the allocation basis used. 2. What was the carrying amount of the building on December 31, 2018? 3. Explain the effect of depreciation on cash flows. e
Answer #1

As asked for answering requirement 1:

Machine A Machine B Machine C
Cost of the machine 10381.94 28888.89 25729.17
Installation costs 800 1100 1100
Renovation costs prior to use 600 1400 1600
Total cost 11781.94 31388.89 28429.17

Total cost of the machine including transportation charges = 62000 +3000 = 65000

The purchase cost of the asset is allocated on the basis of the appraisal value of the assets since the assets are bought estimating the future use based on current asset value.


Allocation of costs to machine A = 65000*11500/(11500+32000+28500) = 10381.94


Allocation of costs to machine B = 65000*32000/(11500+32000+28500) = 28888.89


Allocation of costs to machine A = 65000*28500/(11500+32000+28500) = 25729.17